# Accelerating Mobile Language Model Generation via Hybrid Context and Hardware Coordination

### v1

@@ 4.2. Overall Performance
Per-Token Latency. CoordGenis faster and achieves improvements ranging 1.06–3.81×, as shown in Figure 10.
@@
$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), CoordGen consistently reduces per-token latency by 1.14–3.81× and 1.06–1.78×. These performance gains stem from both the NPU-optimized SD method and the specialized compute graphs. Larger improvements are observed on small-to-medium models (e.g., Qwen2.5 0.5B and 1.5B) as their inference is memory-bound which is preferred by SD. Comparing different datasets, we find that tasks with higher context similarity such as summary, exhibit higher speed gains (1.41–3.80×, 1.24–1.61×) since SD is more effective.
$\bullet$ Compared to baselines with SD, CoordGen consistently improves by 1.11–2.53× than SAM, 1.09–1.80× than EAGLE. This arises because SAM ignores the ineffectiveness of computing short drafts on NPUs while EAGLE suffers from the latency of running another parametric model. Improvements of applying SD on CPU vanilla are moderate (0.92–1.13× for SAM, 0.56–1.18× for EAGLE), as LLM inference on mobile CPU is compute-bound due to the limited CPU capacity.
Energy Consumption. CoordGen is more energy-efficient and reduces energy consumption by 1.07–4.71× (Figure 11).
$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), CoordGen significantly reduces energy consumption by 1.35–4.18× and 1.11–2.50×. The improvements stem jointly from the higher generation speed and the avoidance of CPU computation, allowing energy-efficient utilization of NPUs. On larger models (e.g., Llama3.2 3B), the effectiveness becomes less pronounced (1.42–1.85×) as relatively longer inference time results in the device continues working under a high power consumption.
$\bullet$ Compared to baselines with SD, CoordGen consumes significantly less energy than SAM (1.07–3.07×) and EAGLE (1.77–4.71×). Energy saving of SAM is limited due to frequent verification incurred by rejected drafts. EAGLE suffers from additional computation required by drafter, amplifying energy overhead. For instance, for Qwen2.5 0.5B, CoordGen reduces energy ranging 1.40–3.50× compared with NPU vanilla, while SAM reduces energy by only 1.17–1.54× and EAGLE even increases energy consumption.
@@ 4.3. Memory Overhead
As shown in Figure 12, CoordGen maintains a memory footprint close to NPU vanilla with an additional overhead less than 500 MB across all evaluations. For example, CoordGen requires 3.14 GB for Qwen2.5 1.5B, nearly identical to NPU vanilla (3.05–3.12 GB), while CPU vanilla requires over 6 GB. This gap comes from maintaining separate weights for CPU and NPU in heterogeneous baselines.
The memory overhead of CoordGen primarily arises from distribution calibration and context-relevant database structures, which is manageable for mobile devices with limited memory. In contrast, EAGLE’s memory usage increases noticeably (0.84–1.22 GB) due to maintaining the drafter model, highlighting that CoordGen delivers superior efficiency without sacrificing memory scalability.
@@ 4.4. Latency Breakdown
To analyze the runtime characteristics and latency overhead of CoordGen, we decompose the end-to-end inference latency into prefill, decoding and potential overhead. Measurements are conducted on Oneplus 13 under the summary benchmark, with results shown in Figure 13.
@@
The results show that CoordGen significantly accelerates both prefill and decoding compared to NPU vanilla baseline, with little additional overhead. Specifically, decoding of NPU vanilla baseline consumes a substantial portion of total latency (67–74%), highlighting the significance of introducing specialized optimization. CoordGen directly addresses this by integrating retrieval-based speculative decoding, reducing decoding latency by 5.25–8.15×.
@@ 5. Related Work
Mobile LLM inference optimization. The constrained resources of mobile devices require reducing computation to enable efficient on-device LLM inference (Xu et al., 2024; Subramanian et al., 2025; Ma et al., 2024; Thawakar et al., 2024; Xu et al., 2025a). Several approaches exploit different forms of sparsity. Mixture-of-experts leverages activation sparsity by computing only neurons important to the final output (Yi et al., 2025; Chen et al., 2025c; Wu et al., 2025b; Frantar and Alistarh, 2024). Early-exit reduces layer-level computation by skipping non-critical layers (Venkatesha et al., 2025; Fan et al., 2024). Model compression reduces model size by retaining only the most valuable layers (Yu et al., 2024; Smith et al., 2025). Token pruning eliminates redundant tokens to shorten the context (Liu et al., 2025b; Ma et al., 2023). CoordGen is compatible with these techniques, as it requires no architectural modifications to models or inputs.
Mobile NPU offloading. Modern mobile SoCs are increasingly equipped with high-performance NPUs (OPPO, 2024; Samsung, 2024; Huawei, 2024), offering new opportunities for LLM acceleration (Xue et al., 2024; Chen et al., 2025a; Lu et al., 2025; Xu et al., 2025b; Yin et al., 2025; Wei et al., 2025). PowerInfer-V2 (Xue et al., 2024) integrates parameter swapping with NPU offloading to scale inference to larger LLMs. HeteroLLM (Chen et al., 2025a) introduces tensor partitioning strategies tailored to mobile GPUs and NPUs for efficient parallelization. While these studies primarily focus on accelerating prefill, CoordGen complements them by incorporating speculative decoding to optimize both prefill and decoding.
@@
Speculative decoding uses LLM as a verifier to process multiple draft tokens in parallel, shifting memory-bound decoding to compute-bound (Stern et al., 2018; Chen et al., 2023; Leviathan et al., 2023b). The main difference between these studies is how drafts are generated. Model-based methods employ an auxiliary parametric model, typically a smaller LLM (Li et al., 2024c, b; Cai et al., 2024; Du et al., 2024; Li et al., 2025b). Although these approaches improve the acceptance ratio, they introduce significant drafting overhead (Hu et al., 2024). Retrieval-based methods avoid this cost by using lightweight retrievers for drafting, but often suffer from low acceptance ratios (Saxena, 2023; Somasundaram et al., 2024; Hu et al., 2024; Chen et al., 2025b; Li et al., 2025a; Gritta et al., 2025). CoordGen advances R-SD through hardware–algorithm co-design, integrating it with mobile NPUs for efficient acceleration.
@@ 6. Conclusions
In this paper, we propose CoordGen, the first framework designed to align retrieval-based speculative decoding with mobile NPUs to optimize context-augmented generation. CoordGen incorporates novel techniques such as progressive graph scheduling and in-context calibration to address the gaps between speculative decoding and NPUs. Extensive experiments demonstrate CoordGen achieves outstanding and consistent performance gains.

### v3

@@ 4.2. Overall Performance
Per-Token Latency. sd.npuis faster and achieves improvements ranging 1.06–3.81×, as shown in Figure 10.
@@
$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), sd.npu consistently reduces per-token latency by 1.14–3.81× and 1.06–1.78×. These performance gains stem from both the NPU-optimized SD method and the specialized compute graphs. Larger improvements are observed on small-to-medium models (e.g., Qwen2.5 0.5B and 1.5B) as their inference is memory-bound which is preferred by SD. Comparing different datasets, we find that tasks with higher context similarity such as summary, exhibit higher speed gains (1.41–3.80×, 1.24–1.61×) since SD is more effective.
$\bullet$ Compared to baselines with SD, sd.npu consistently improves by 1.11–2.53× than SAM, 1.09–1.80× than EAGLE. This arises because SAM ignores the ineffectiveness of computing short drafts on NPUs while EAGLE suffers from the latency of running another parametric model. Improvements of applying SD on CPU vanilla are moderate (0.92–1.13× for SAM, 0.56–1.18× for EAGLE), as LLM inference on mobile CPU is compute-bound due to the limited CPU capacity.
Energy Consumption. sd.npu is more energy-efficient and reduces energy consumption by 1.07–4.71× (Figure 11).
$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), sd.npu significantly reduces energy consumption by 1.35–4.18× and 1.11–2.50×. The improvements stem jointly from the higher generation speed and the avoidance of CPU computation, allowing energy-efficient utilization of NPUs. On larger models (e.g., Llama3.2 3B), the effectiveness becomes less pronounced (1.42–1.85×) as relatively longer inference time results in the device continues working under a high power consumption.
$\bullet$ Compared to baselines with SD, sd.npu consumes significantly less energy than SAM (1.07–3.07×) and EAGLE (1.77–4.71×). Energy saving of SAM is limited due to frequent verification incurred by rejected drafts. EAGLE suffers from additional computation required by drafter, amplifying energy overhead. For instance, for Qwen2.5 0.5B, sd.npu reduces energy ranging 1.40–3.50× compared with NPU vanilla, while SAM reduces energy by only 1.17–1.54× and EAGLE even increases energy consumption.
@@ 4.3. Memory Overhead
As shown in Figure 12, sd.npu maintains a memory footprint close to NPU vanilla with an additional overhead less than 500 MB across all evaluations. For example, sd.npu requires 3.14 GB for Qwen2.5 1.5B, nearly identical to NPU vanilla (3.05–3.12 GB), while CPU vanilla requires over 6 GB. This gap comes from maintaining separate weights for CPU and NPU in heterogeneous baselines.
The memory overhead of sd.npu primarily arises from distribution calibration and context-relevant database structures, which is manageable for mobile devices with limited memory. In contrast, EAGLE’s memory usage increases noticeably (0.84–1.22 GB) due to maintaining the drafter model, highlighting that sd.npu delivers superior efficiency without sacrificing memory scalability.
@@ 4.4. Latency Breakdown
To analyze the runtime characteristics and latency overhead of sd.npu, we decompose the end-to-end inference latency into prefill, decoding and potential overhead. Measurements are conducted on Oneplus 13 under the summary benchmark, with results shown in Figure 13.
@@
The results show that sd.npu significantly accelerates both prefill and decoding compared to NPU vanilla baseline, with little additional overhead. Specifically, decoding of NPU vanilla baseline consumes a substantial portion of total latency (67–74%), highlighting the significance of introducing specialized optimization. sd.npu directly addresses this by integrating retrieval-based speculative decoding, reducing decoding latency by 5.25–8.15×.
@@ 5. Related Work
Mobile LLM inference optimization. The constrained resources of mobile devices require reducing computation to enable efficient on-device LLM inference (Xu et al., 2024; Subramanian et al., 2025; Ma et al., 2024; Thawakar et al., 2024; Xu et al., 2025a). Several approaches exploit different forms of sparsity. Mixture-of-experts leverages activation sparsity by computing only neurons important to the final output (Yi et al., 2025; Chen et al., 2025c; Wu et al., 2025b; Frantar and Alistarh, 2024). Early-exit reduces layer-level computation by skipping non-critical layers (Venkatesha et al., 2025; Fan et al., 2024). Model compression reduces model size by retaining only the most valuable layers (Yu et al., 2024; Smith et al., 2025). Token pruning eliminates redundant tokens to shorten the context (Liu et al., 2025b; Ma et al., 2023). sd.npu is compatible with these techniques, as it requires no architectural modifications to models or inputs.
Mobile NPU offloading. Modern mobile SoCs are increasingly equipped with high-performance NPUs (OPPO, 2024; Samsung, 2024; Huawei, 2024), offering new opportunities for LLM acceleration (Xue et al., 2024; Chen et al., 2025a; Lu et al., 2025; Xu et al., 2025b; Yin et al., 2025; Wei et al., 2025). PowerInfer-V2 (Xue et al., 2024) integrates parameter swapping with NPU offloading to scale inference to larger LLMs. HeteroLLM (Chen et al., 2025a) introduces tensor partitioning strategies tailored to mobile GPUs and NPUs for efficient parallelization. While these studies primarily focus on accelerating prefill, sd.npu complements them by incorporating speculative decoding to optimize both prefill and decoding.
@@
Speculative decoding uses LLM as a verifier to process multiple draft tokens in parallel, shifting memory-bound decoding to compute-bound (Stern et al., 2018; Chen et al., 2023; Leviathan et al., 2023b). The main difference between these studies is how drafts are generated. Model-based methods employ an auxiliary parametric model, typically a smaller LLM (Li et al., 2024c, b; Cai et al., 2024; Du et al., 2024; Li et al., 2025b). Although these approaches improve the acceptance ratio, they introduce significant drafting overhead (Hu et al., 2024). Retrieval-based methods avoid this cost by using lightweight retrievers for drafting, but often suffer from low acceptance ratios (Saxena, 2023; Somasundaram et al., 2024; Hu et al., 2024; Chen et al., 2025b; Li et al., 2025a; Gritta et al., 2025). sd.npu advances R-SD through hardware–algorithm co-design, integrating it with mobile NPUs for efficient acceleration.
@@ 6. Conclusions
In this paper, we propose sd.npu, the first framework designed to align retrieval-based speculative decoding with mobile NPUs to optimize context-augmented generation. sd.npu incorporates novel techniques such as progressive graph scheduling and in-context calibration to address the gaps between speculative decoding and NPUs. Extensive experiments demonstrate sd.npu achieves outstanding and consistent performance gains.